OpenStudy (yaya090600):

Derive the equation of the parabola with a focus at (−5, 5) and a directrix of y = −1.

OpenStudy (yaya090600):

f(x) = −one twelfth (x − 5)2 + 2 f(x) = one twelfth (x − 5)2 + 2 f(x) = −one twelfth (x + 5)2 + 2 f(x) = one twelfth (x + 5)2 + 2

OpenStudy (yaya090600):

@StudyGurl14 @mjmahmood

OpenStudy (anonymous):

yes

OpenStudy (yaya090600):

can you help?

OpenStudy (anonymous):

The equation of a parabola with focus in (h,k+p) and directrix y=k-p, is, (x-h)^2=4p(y-k)In the first case, h=-5\\ k+p=-5\\y=k-p=7Combining the last two relations, k=1,p=-6. So the equation of the parabola is, (x+5)^2=-24(y-1)

OpenStudy (anonymous):

did that help

OpenStudy (yaya090600):

so its the last one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

was it correct

OpenStudy (yaya090600):

i didnt check it yet

OpenStudy (yaya090600):

i still need help with 2 more questions

OpenStudy (anonymous):

i can help you

OpenStudy (yaya090600):

Derive the equation of the parabola with a focus at (2, −1) and a directrix of y = −one half. f(x) = −(x + 2)2 − three fourths f(x) = (x + 2)2 + three fourths f(x) = −(x − 2)2 + three fourths f(x) = −(x − 2)2 − three fourths

OpenStudy (yaya090600):

Derive the equation of the parabola with a focus at (−2, 4) and a directrix of y = 6. Put the equation in standard form. f(x) = one fourthx2 − x + 4 f(x) = −one fourthx2 − x + 4 f(x) = one fourthx2 − x + 5 f(x) = −one fourthx2 − x + 5

OpenStudy (anonymous):

so -4(y - 5) = (x - -2)²

OpenStudy (anonymous):

The first thing to do is to decide which way up this is and whether its a side ways one or not. Clearly there are 4 different types open up or down or open to the left or right. The focus is inside the parabola and the directrix is a line on the outside but the same distance from the vertex as the vertex is from the focus. Because the focus is below the directrix then this is a parabola that opens downwards and has an equation that has an x² 4p(y - k) = (x - h)² is the general form of and opening up or down one p is the distance from the directrix to the vertex or twice the distance from the directrix to the focus (h,k) is the co-ords of the vertex 2|p| = 2 so |p| = 1 (the focus is at y = 4 and the directrix is y = 6) difference 2 but because we are facing downwards p = -1 Vertex is at (-2,5) the y co-ord is halfway between the focus and the directrix The x bit is -2 for the focus and the vertex so the equation becomes so -4(y - 5) = (x - -2)² -4(y - 5) = (x + 2)² -4y +20 = x² +4x + 4 -4y = x² +4x - 16 4y = -x² -4x + 16 y = -1/4x² -x + 4

OpenStudy (yaya090600):

wait whats the first one?

OpenStudy (anonymous):

what do you mean the first one

OpenStudy (anonymous):

o i see what you mean thats how to do it

OpenStudy (yaya090600):

I dont know which one is the answer tho

OpenStudy (yaya090600):

its either a or b right? @mjmahmood

OpenStudy (yaya090600):

its d lol

OpenStudy (yaya090600):

i got 100% thanks!

OpenStudy (anonymous):

ok good